# The Potential and Challenges of Evaluating Attitudes, Opinions, and Values in Large Language Models

### v2

@@ A.2 Models Deployed in the Surveyed Works
The distribution of these 37 models is illustrated in Figure 3. From the figure, we can observe that the closed-source GPT models are the most popular, with GPT-3.5 being the most frequently deployed model with 27 instances, followed by GPT-3 with 21 instances, and GPT-4 with 18 instances. The open-source models like Llama-2 and Mistral Jiang et al. (2023a) also have notable counts, with 15 and 8 instances respectively. However, a good amount of models such as Codex Chen et al. (2021) and MPT Team (2023) are among the least frequently deployed, each appearing only once.
@@ A.3 Task Design
We show in this section a brief introduction to the task design for querying LLMs for AOV with a few simple examples. Most works use original surveys or questionnaires designed for human participants, which are mostly closed-ended (e.g. Argyle et al., 2023; Santurkar et al., 2023; Hwang et al., 2023; Wang et al., 2024b), for querying the LLMs. Figure 4 and 5 showcase the close-ended questions without or with appended persona input prompt, respectively. Some focus on open-ended settings to emphasize textual output (e.g. Jiang et al., 2022a; Simmons, 2023; Benkler et al., 2023). Figure 6 presents a prompt template asking for opinions in an open-ended setting. Röttger et al. (2024) compare closed-ended and open-ended settings with further splitting the open-ended setting into a “forced” open-ended setting by adding a sentence, “Take a clear stance”, and a “fully unconstrained” open-ended setting, to test model robustness, as shown in Figure 7. These settings are further employed by Wright et al. (2024).
@@ D. Should not be done
Non-US-Centric Public Opinion Polls von der Heyde et al. (2023); Kalinin (2023); Durmus et al. (2023)
Non Public Opinion Polls Jiang et al. (2022a); Rozado (2023); Rozado (2024); Rosenbusch et al. (2023); Wu et al. (2023); Hartmann et al. (2023); Chalkidis and Brandl (2024); Haller et al. (2024); Feng et al. (2023); Röttger et al. (2024); Ceron et al. (2024); Bang et al. (2024); Wright et al. (2024)
Values Value Orientation of LLMs Simmons (2023); Benkler et al. (2023); Fraser et al. (2022); Cao et al. (2023); Arora et al. (2023); Johnson et al. (2022); Abdulhai et al. (2023); Tanmay et al. (2023); Haemmerl et al. (2023); Talat et al. (2022)
Datasets and Frameworks Benkler et al. (2022); Jin et al. (2022); Sorensen et al. (2024); Klingefjord et al. (2024); Rao et al. (2023); Agarwal et al. (2024); Hendrycks et al. (2023); Scherrer et al. (2024); Ren et al. (2024); Aharoni et al. (2024); Bonagiri et al. (2024)
Others Persona and Personality Miotto et al. (2022); Kovač et al. (2023); Caron and Srivastava (2023); Cheng et al. (2023a); Cheng et al. (2023b); Jiang et al. (2024); Shu et al. (2024); Hu and Collier (2024)
Theory-of-Mind Sap et al. (2022); Li et al. (2023b); Kosinski (2024)
@@
Table 1: Overview of related works for studying AOV in LLMs.

### v3

@@ A.2 Models Deployed in the Surveyed Works
The distribution of these 37 models is illustrated in Figure 3. From the figure, we can observe that the closed-source GPT models are the most popular, with GPT-3.5 being the most frequently deployed model with 30 instances, followed by GPT-3 with 22 instances, and GPT-4 with 21 instances. The open-source models like Llama-2 and Mistral Jiang et al. (2023a) also have notable counts, with 17 and 9 instances respectively. However, a good amount of models such as Codex Chen et al. (2021) and MPT Team (2023) are among the least frequently deployed, each appearing only once.
@@ A.3 Task Design
We show in this section a brief introduction to the task design for querying LLMs for AOVs with a few simple examples. Most works use original surveys or questionnaires designed for human participants, which are mostly closed-ended (e.g. Argyle et al., 2023; Santurkar et al., 2023; Hwang et al., 2023; Wang et al., 2024b), for querying the LLMs. Figure 4 and 5 showcase the close-ended questions without or with appended persona input prompt, respectively. Some focus on open-ended settings to emphasize textual output (e.g. Jiang et al., 2022a; Simmons, 2023; Benkler et al., 2023). Figure 6 presents a prompt template asking for opinions in an open-ended setting. Röttger et al. (2024) compare closed-ended and open-ended settings with further splitting the open-ended setting into a “forced” open-ended setting by adding a sentence, “Take a clear stance”, and a “fully unconstrained” open-ended setting, to test model robustness, as shown in Figure 7. These settings are further employed by Wright et al. (2024).
@@ D. Should not be done
Non-US-Centric Public Opinion Polls von der Heyde et al. (2023); Kalinin (2023); Geng et al. (2024); Durmus et al. (2024)
Additional Data Source Used Jiang et al. (2022a); Rozado (2023); Rozado (2024); Rosenbusch et al. (2023); Wu et al. (2023); Hartmann et al. (2023); Chalkidis and Brandl (2024); Haller et al. (2024); Feng et al. (2023); Röttger et al. (2024); Ceron et al. (2024); Bang et al. (2024); Wright et al. (2024)
Values Value Orientation of LLMs Simmons (2023); Benkler et al. (2023); Fraser et al. (2022); Cao et al. (2023); Arora et al. (2023); Johnson et al. (2022); Abdulhai et al. (2023); Tanmay et al. (2023); Haemmerl et al. (2023); Talat et al. (2022); Jinnai (2024)
Datasets and Frameworks Benkler et al. (2022); Jin et al. (2022); Sorensen et al. (2024); Klingefjord et al. (2024); Rao et al. (2023a); Agarwal et al. (2024); Hendrycks et al. (2023); Scherrer et al. (2023); Ren et al. (2024); Aharoni et al. (2024); Bonagiri et al. (2024); Yu et al. (2024)
Others Persona and Personality Miotto et al. (2022); Kovač et al. (2023); Caron and Srivastava (2023); Cheng et al. (2023a); Cheng et al. (2023b); Rao et al. (2023b); Jiang et al. (2024); Shu et al. (2024); Hu and Collier (2024)
Theory-of-Mind Sap et al. (2022); Li et al. (2023b); Kosinski (2024); Strachan et al. (2024)
@@
Table 1: Overview of related works for studying AOVs in LLMs.